You hear on the news that the​ S&P 500 was down 2.2% today relative to the​ risk-free rate​ (the market's excess return was −2.2%).

You are thinking about your portfolio and your investments in Zynga and Proctor and Gamble.

a. If​ Zynga's beta is 1.1​, what is your best guess as to​ Zynga's excess return​ today?

b. If Proctor and​ Gamble's beta is 0.6​, what is your best guess as to​ P&G's excess return​ today?

a. If​ Zynga's beta is 1.1​, what is your best guess as to​ Zynga's excess return​ today? ​Zynga's excess return today is? (Round to one decimal​ place.)
Answer #1

Beta = Asset's Excess Return / Market's Excess Return

Asset's Excess Return = Beta * Market's Excess Return

a). Zynga's Excess Return = 1.1 * (-2.2%) = -2.4%

b). Proctor and Gamble's Excess Return = 0.6 * (-2.2%) = -1.3%